KingCrazy:

Write the set of integers les than -6 or greater than 21 in set-builder notation. Write the interval in interval notation

jhonyy9:

the set of integers les than -6 or greater than 21 did you wrote it correct : ... < -6 or 21 < ... ?

surjithayer:

|x-a|>b means x-a<-b,x-a>b or x<a-b,x>a+b a-b=-6 a+b=21 add 2a=15 a=15/2=7.5 b=21-7.5=13.5 so |x-7.5|>13.5 or \[(-\infty,-6)U(21,\infty)\]